# Bangkok 10330, Thailand

### v1

@@ 3.2 Dominant contributions to non-gaussianity and its covariant definition
$f_{\text{NL}}\equiv\frac{5}{12}\frac{\mathscr{A}^{*}}{k_{1}^{3}+k_{2}^{3}+k_{3}^{3}}\,.$ (28)
@@
At LO $f_{\text{NL}}$ is a function of ratios of momenta, while at NLO it acquires a logarithmic dependence, as seen in (19) with $\tau$ replaced by $\tau^{*}$. Note that $-\tau^{*}k$ is the physical momentum measured at time $\tau^{*}$ in units of the inflation scale $M_{\text{I}}\simeq H$: $-\tau^{*}k=k_{\text{phys}}/M_{\text{I}}=2\pi/(\lambda M_{\text{I}})$, with $\lambda$ some physical distance probed by CMB observations in the range $\sim 0.01-100$ Mpc, when the Universe was about 1000 times smaller than today. It follows that for momenta in this range, the dominant contribution to NLO relevant for observations comes from the logarithm $\ln(-\tau^{*}k_{\text{t}})$ in the first two lines of (19), as well as when a ratio of momenta becomes large near the squeezed limit that we study in the next section.
Using (18) and (19) in (28), one then finds Antoniadis:2024abm :
@@
$\begin{split}f_{\text{NL}}\simeq\frac{5}{12}&\bigg{\{}\varepsilon_{1}\widetilde{K}_{1}+\varepsilon_{2}+\left[-\varepsilon_{1}(\varepsilon_{2}+4\varepsilon_{1})\widetilde{K}_{1}+\varepsilon_{2}(\varepsilon_{3}-2\varepsilon_{2}-4\varepsilon_{1})\right]\ln(-\tau k_{\text{t}})\\ &\qquad\qquad\quad\,-\,(\varepsilon_{2}+2\varepsilon_{1})^{2}\left(\ln{8k_{1}k_{2}k_{3}\over k_{\text{t}}^{3}}\right)_{\text{sq}}\bigg{\}},\end{split}$ (29)
@@
where we defined $\widetilde{K}_{1}(k_{1},k_{2},k_{3})=K_{1}(k_{1},k_{2},k_{3})/K_{2}(k_{1},k_{2},k_{3})$ [^6] Note that this $\widetilde{K}_{1}$ is different from the definition in Antoniadis:2024abm by the numerical factor $5/12$. and the term in the second line contributes only in the squeezed limit where one of the momenta is small, while we also dropped for the moment the star on $\tau$ for notational simplicity.
@@
Note however that the definition (28) is not time independent since the proportionality factor between the bispectrum $\mathcal{B}$ and the 3-point amplitude $\mathscr{A}$ in (16) acquires time dependence at the next order. Actually, using the lowest order expression of the power spectrum $\mathcal{P}_{k}^{\text{LO}}=H^{2}/(4k^{3}\varepsilon_{1})$ and (26), one can rewrite (28) as
$f_{\text{NL}}={5\over 6}{\mathcal{B}\over\mathcal{P}_{k_{1}}^{\text{LO}}\mathcal{P}_{k_{2}}^{\text{LO}}+\mathcal{P}_{k_{2}}^{\text{LO}}\mathcal{P}_{k_{3}}^{\text{LO}}+\mathcal{P}_{k_{3}}^{\text{LO}}\mathcal{P}_{k_{1}}^{\text{LO}}},$ (30)
@@
which is manifestly time dependent at NLO and beyond. A time independent definition would therefore be obtained by replacing in the denominator the LO power spectrum by its full conserved (time independent) expression Burrage:2011hd :
$\widehat{f}_{\text{NL}}={5\over 6}{\mathcal{B}(k_{1},k_{2},k_{3})\over\mathcal{P}_{k_{1}}\mathcal{P}_{k_{2}}+\mathcal{P}_{k_{2}}\mathcal{P}_{k_{3}}+\mathcal{P}_{k_{3}}\mathcal{P}_{k_{1}}}.$ (31)
@@
Similarly to the bispectrum, the dominant contribution to the power spectrum up to NLO contains logarithmic corrections and reads [^7] This is independent of $\tau$ as demonstrated in Appendix D of Antoniadis:2024abm and also in Section 4.3. Antoniadis:2024abm :
$\mathcal{P}_{k}\simeq\frac{H^{2}}{4k^{3}\varepsilon_{1}}\left[1-(\varepsilon_{2}+2\varepsilon_{1})\ln(-k\tau)\right].$ (32)
@@
It follows that $\widehat{f}_{\text{NL}}$ at NLO contains an extra term compared to the expression of $f_{\text{NL}}$ in (29), given by
@@
$\begin{split}\widehat{f}_{\text{NL,NLO}}&\simeq f_{\text{NL,NLO}}\\ &\quad+(\varepsilon_{2}+2\varepsilon_{1})f_{\text{NL,LO}}{k_{3}^{3}\ln(k_{1}k_{2}\tau^{2})+k_{2}^{3}\ln(k_{3}k_{1}\tau^{2})+k_{1}^{3}\ln(k_{2}k_{3}\tau^{2})\over k_{1}^{3}+k_{2}^{3}+k_{3}^{3}},\end{split}$ (33)
@@
where the symbol $\simeq$ denotes the dominant (logarithmic) contribution, as in (29) and (32). In order to separate the two dominant contributions as in (29), we multiply the arguments of the logarithms by $k_{\text{t}}^{2}/k_{\text{t}}^{2}$ and separate the universal part $\ln(k_{\text{t}}\tau)^{2}$ from those which are enhanced in the squeezed limits, to obtain:
@@
$\qquad\quad\quad-(\varepsilon_{2}+2\varepsilon_{1})^{2}\left(\ln{8k_{1}k_{2}k_{3}\over k_{\text{t}}^{3}}\right)_{\text{sq}}+(\varepsilon_{2}+2\varepsilon_{1})\left((\varepsilon_{1}\widetilde{K}_{1}+\varepsilon_{2})\ln{8k_{1}k_{2}k_{3}\over k_{\text{t}}^{3}}\right)_{\text{sq}},$ (34)
@@
$\widehat{f}_{\text{NL}}\simeq\frac{5}{12}\left\{\varepsilon_{1}\widetilde{K}_{1}+\varepsilon_{2}+\varepsilon_{2}(\varepsilon_{1}\widetilde{K}_{1}+\varepsilon_{3})\ln(-\tau k_{\text{t}})\right\}.$ (35)
It is now straightforward to show its time independence up to NLO by using the expansion formula (23). Note that the squeezed limit divergence has been cancelled in the covariant version $\widehat{f}_{\text{NL}}$ of the parameter $f_{\text{NL}}$.
@@ 4.1 Squeezed limit of the bispectrum and power spectrum
Let us evaluate the squeezed limit $k_{1}\to 0$ of the bispectrum up to NLO. We start with the bispectrum around the pivot scale $\tau^{*}$ (26). Since $k_{1}\sim 0$, the other momenta have an almost equal size, $k\equiv k_{2}=k_{3}+\mathcal{O}(k_{1})$. Using $K_{1}(k_{1},k_{2},k_{3})\simeq 4k^{3}$ and $K_{2}(k_{1},k_{2},k_{3})\simeq 2k^{3}$ in this limit, we obtain
$\mathcal{B}(k_{1},k_{2},k_{3})_{k_{1}\to 0}=\frac{H^{*4}}{32k_{1}^{3}k^{6}\varepsilon_{1}^{*2}}\mathscr{A}^{*}_{k_{1}\rightarrow 0},$ (36)
$\begin{split}&\mathscr{A}^{*}_{k_{1}\rightarrow 0}=2k^{3}\Bigg{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}+2\varepsilon_{1}^{*2}+5\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\\ &\qquad\qquad\qquad~{}-\big{(}8\varepsilon_{1}^{*2}+6\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\big{)}\big{[}\gamma_{E}-1+\ln(-2k\tau^{*})\big{]}\\ &\qquad\qquad\qquad~{}-\big{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}\big{)}^{2}\ln\left(\frac{k_{1}}{k}\right)\Bigg{)}\left[1+\mathcal{O}\left(\frac{k_{1}}{k},\varepsilon^{*3},\tau\right)\right].\end{split}$ (37)
Let us proceed to the power spectrum. As done for the bispectrum, it is convenient to introduce the amplitude $\widehat{\mathcal{P}}_{k}$ of the power spectrum as
$\mathcal{P}_{k}=\frac{H^{*2}}{4k^{3}\varepsilon_{1}^{*}}\mathscr{P}^{*}_{k},$ (38)
where we adopted the pivot scale $\tau^{*}$. We have seen that the amplitude of the bispectrum $\mathscr{A}^{*}$ at NLO is of second order in $\varepsilon_{i}^{*}$. In order to reproduce this from the power spectrum, we need $\mathscr{P}^{*}_{k}$ up to second order in $\varepsilon_{i}^{*}$, which is at the next-to-next-to-leading order (N2LO) for the power spectrum because $\mathscr{P}^{*}_{k}$ at LO is 1. In the late-time limit, the amplitude $\mathscr{P}^{*}_{k}$ up to N2LO with any pivot scale $\tau^{*}$ is given by Gong:2001he ; Schwarz:2001vv ; Auclair:2022yxs ; Bianchi:2024qyp ; Antoniadis:2024abm
$\mathscr{P}^{*}_{k}=\mathcal{A}+\mathcal{B}\ln(-k\tau^{*})+\mathcal{C}\big{(}\ln(-k\tau^{*})\big{)}^{2}+\mathcal{O}\big{(}\varepsilon^{*3},\tau\big{)},$ (39)
@@
$\begin{split}&\mathcal{A}=1-2(\lambda+1)\varepsilon_{1}^{*}-\lambda\varepsilon_{2}^{*}+\frac{1}{2}\big{(}\pi^{2}+4\lambda^{2}+4\lambda-6\big{)}\varepsilon_{2}^{*2}\\ &\quad+\frac{1}{12}\big{(}7\pi^{2}+12\lambda^{2}-12\lambda-72\big{)}\varepsilon_{1}^{*}\varepsilon_{2}^{*}+\frac{1}{8}\big{(}\pi^{2}+4\lambda^{2}-8\big{)}\varepsilon_{2}^{*2}+\frac{1}{24}\big{(}\pi^{2}-12\lambda^{2}\big{)}\varepsilon_{2}^{*}\varepsilon_{3}^{*},\end{split}$ (40)
$\begin{split}&\mathcal{B}=-2\varepsilon_{1}^{*}-\varepsilon_{2}^{*}+(4\lambda+2)\varepsilon_{1}^{*2}+(2\lambda-1)\varepsilon_{1}^{*}\varepsilon_{2}^{*}+\lambda\big{(}\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\big{)},\end{split}$ (41)
$\begin{split}&\mathcal{C}=2\varepsilon_{1}^{*2}+\varepsilon_{1}^{*}\varepsilon_{2}^{*}+\frac{1}{2}\varepsilon_{2}^{*2}-\frac{1}{2}\varepsilon_{2}^{*}\varepsilon_{3}^{*},\end{split}$ (42)
@@ 4.2 Consistency relation at NLO
Finally, computing the right hand side of (1) by using (38) with (39), we obtain
$\begin{split}-\mathcal{P}_{k_{1}}\left(3+k_{2}\frac{d}{dk_{2}}\right)\mathcal{P}_{k_{2}}&=\frac{H^{*4}}{16k_{1}^{3}k_{2}^{3}\varepsilon_{1}^{*2}}\Bigg{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}+2\varepsilon_{1}^{*2}+5\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\\ &\qquad~{}-\big{(}8\varepsilon_{1}^{*2}+6\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\big{)}\big{[}\gamma_{E}-1+\ln(-2k_{2}\tau^{*})\big{]}\\ &\qquad~{}-\big{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}\big{)}^{2}\ln\left(\frac{k_{1}}{k_{2}}\right)\Bigg{)}+\mathcal{O}\big{(}\varepsilon^{*3},\tau\big{)},\end{split}$ (43)
which, under $k=k_{2}=k_{3}+\mathcal{O}(k_{1})$, is equal to the bispectrum in the squeezed limit (36) with (37), thereby verifying the consistency relation at NLO. Note that this has been verified for any choice of the pivot scale $\tau^{*}$.

### v2

@@ 3.2 Dominant contributions to non-gaussianity and its conserved definition
$f_{\text{NL}}\equiv\frac{5}{12}\frac{\mathscr{A}^{*}}{k_{1}^{3}+k_{2}^{3}+k_{3}^{3}}\,.$ (3.13)
@@
At LO $f_{\text{NL}}$ is a function of ratios of momenta, while at NLO it acquires a logarithmic dependence, as seen in (3.4) with $\tau$ replaced by $\tau^{*}$. Note that $-\tau^{*}k$ is the physical momentum measured at time $\tau^{*}$ in units of the inflation scale $M_{\text{I}}\simeq H$: $-\tau^{*}k=k_{\text{phys}}/M_{\text{I}}=2\pi/(\lambda M_{\text{I}})$, with $\lambda$ some physical distance probed by CMB observations in the range $\sim 0.01-100$ Mpc, when the Universe was about 1000 times smaller than today. It follows that for momenta in this range, the dominant contribution to NLO relevant for observations comes from the logarithm $\ln(-\tau^{*}k_{\text{t}})$ in the first two lines of (3.4), as well as when a ratio of momenta becomes large near the squeezed limit that we study in the next section.
Using (3.3) and (3.4) in (3.13), one then finds [4]:
@@
$\begin{split}f_{\text{NL}}\simeq\frac{5}{12}&\bigg{\{}\varepsilon_{1}\widetilde{K}_{1}+\varepsilon_{2}+\left[-\varepsilon_{1}(\varepsilon_{2}+4\varepsilon_{1})\widetilde{K}_{1}+\varepsilon_{2}(\varepsilon_{3}-2\varepsilon_{2}-4\varepsilon_{1})\right]\ln(-\tau k_{\text{t}})\\ &\qquad\qquad\quad\,-\,(\varepsilon_{2}+2\varepsilon_{1})^{2}\left(\ln{\frac{8k_{1}k_{2}k_{3}}{k_{\text{t}}^{3}}}\right)_{\text{sq}}\bigg{\}},\end{split}$ (3.14)
@@
where we defined $\widetilde{K}_{1}(k_{1},k_{2},k_{3})=K_{1}(k_{1},k_{2},k_{3})/K_{2}(k_{1},k_{2},k_{3})$ [^6] Note that this $\widetilde{K}_{1}$ is different from the definition in [4] by the numerical factor $5/12$. and the term in the second line contributes only in the squeezed limit where one of the momenta is small, while we also dropped for the moment the star on $\tau$ for notational simplicity.
@@
Note however that the definition (3.13) is not time independent since the proportionality factor between the bispectrum $\mathcal{B}$ and the 3-point amplitude $\mathscr{A}$ in (3.1) acquires time dependence at the next order. Actually, using the lowest order expression of the power spectrum $\mathcal{P}_{k}^{\text{LO}}=H^{2}/(4k^{3}\varepsilon_{1})$ and (3.11), one can rewrite (3.13) as
$f_{\text{NL}}=\frac{5}{6}\frac{\mathcal{B}}{\mathcal{P}_{k_{1}}^{\text{LO}}\mathcal{P}_{k_{2}}^{\text{LO}}+\mathcal{P}_{k_{2}}^{\text{LO}}\mathcal{P}_{k_{3}}^{\text{LO}}+\mathcal{P}_{k_{3}}^{\text{LO}}\mathcal{P}_{k_{1}}^{\text{LO}}},$ (3.15)
@@
which is manifestly time dependent at NLO and beyond. A time independent definition would therefore be obtained by replacing in the denominator the LO power spectrum by its full conserved (time independent) expression [16]:
$\widehat{f}_{\text{NL}}=\frac{5}{6}\frac{\mathcal{B}(k_{1},k_{2},k_{3})}{\mathcal{P}_{k_{1}}\mathcal{P}_{k_{2}}+\mathcal{P}_{k_{2}}\mathcal{P}_{k_{3}}+\mathcal{P}_{k_{3}}\mathcal{P}_{k_{1}}}.$ (3.16)
@@
Similarly to the bispectrum, the dominant contribution to the power spectrum up to NLO contains logarithmic corrections and reads [^7] This is independent of $\tau$ as demonstrated in Appendix D of [4] and also in Section 4.3. [4]:
$\mathcal{P}_{k}\simeq\frac{H^{2}}{4k^{3}\varepsilon_{1}}\left[1-(\varepsilon_{2}+2\varepsilon_{1})\ln(-k\tau)\right].$ (3.17)
@@
It follows that $\widehat{f}_{\text{NL}}$ at NLO contains an extra term compared to the expression of $f_{\text{NL}}$ in (3.14), given by
@@
$\begin{split}\widehat{f}_{\text{NL,NLO}}&\simeq f_{\text{NL,NLO}}\\ &\quad+(\varepsilon_{2}+2\varepsilon_{1})f_{\text{NL,LO}}\frac{k_{3}^{3}\ln(k_{1}k_{2}\tau^{2})+k_{2}^{3}\ln(k_{3}k_{1}\tau^{2})+k_{1}^{3}\ln(k_{2}k_{3}\tau^{2})}{k_{1}^{3}+k_{2}^{3}+k_{3}^{3}},\end{split}$ (3.18)
@@
where the symbol $\simeq$ denotes the dominant (logarithmic) contribution, as in (3.14) and (3.17). In order to separate the two dominant contributions as in (3.14), we multiply the arguments of the logarithms by $k_{\text{t}}^{2}/k_{\text{t}}^{2}$ and separate the universal part $\ln(k_{\text{t}}\tau)^{2}$ from those which are enhanced in the squeezed limits, to obtain:
@@
$\quad-(\varepsilon_{2}+2\varepsilon_{1})^{2}\left(\ln{\frac{8k_{1}k_{2}k_{3}}{k_{\text{t}}^{3}}}\right)_{\text{sq}}+(\varepsilon_{2}+2\varepsilon_{1})\left((\varepsilon_{1}\widetilde{K}_{1}+\varepsilon_{2})\ln{\frac{8k_{1}k_{2}k_{3}}{k_{\text{t}}^{3}}}\right)_{\text{sq}},$ (3.19)
@@
$\widehat{f}_{\text{NL}}\simeq\frac{5}{12}\left\{\varepsilon_{1}\widetilde{K}_{1}+\varepsilon_{2}+\varepsilon_{2}(\varepsilon_{1}\widetilde{K}_{1}+\varepsilon_{3})\ln(-\tau k_{\text{t}})\right\}.$ (3.20)
It is now straightforward to show its time independence up to NLO by using the expansion formula (3.8). Note that the squeezed limit divergence has been cancelled in the conserved version $\widehat{f}_{\text{NL}}$ of the parameter $f_{\text{NL}}$.
@@ 4.1 Squeezed limit of the bispectrum and power spectrum
Let us evaluate the squeezed limit $k_{1}\to 0$ of the bispectrum up to NLO. We start with the bispectrum around the pivot scale $\tau^{*}$ (3.11). Since $k_{1}\sim 0$, the other momenta have an almost equal size, $k\equiv k_{2}=k_{3}+\mathcal{O}(k_{1})$. Using $K_{1}(k_{1},k_{2},k_{3})\simeq 4k^{3}$ and $K_{2}(k_{1},k_{2},k_{3})\simeq 2k^{3}$ in this limit, we obtain
$\mathcal{B}(k_{1},k_{2},k_{3})_{k_{1}\to 0}=\frac{H^{*4}}{32k_{1}^{3}k^{6}\varepsilon_{1}^{*2}}\mathscr{A}^{*}_{k_{1}\rightarrow 0},$ (4.1)
$\begin{split}&\mathscr{A}^{*}_{k_{1}\rightarrow 0}=2k^{3}\Bigg{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}+2\varepsilon_{1}^{*2}+5\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\\ &\qquad\qquad\qquad~{}-\big{(}8\varepsilon_{1}^{*2}+6\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\big{)}\big{[}\gamma_{E}-1+\ln(-2k\tau^{*})\big{]}\\ &\qquad\qquad\qquad~{}-\big{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}\big{)}^{2}\ln\left(\frac{k_{1}}{k}\right)\Bigg{)}\left[1+\mathcal{O}\left(\frac{k_{1}}{k},\varepsilon^{*3},\tau\right)\right].\end{split}$ (4.2)
Let us proceed to the power spectrum. As done for the bispectrum, it is convenient to introduce the amplitude $\mathscr{P}_{k}$ of the power spectrum as
$\mathcal{P}_{k}=\frac{H^{*2}}{4k^{3}\varepsilon_{1}^{*}}\mathscr{P}^{*}_{k},$ (4.3)
where we adopted the pivot scale $\tau^{*}$. We have seen that the amplitude of the bispectrum $\mathscr{A}^{*}$ at NLO is of second order in $\varepsilon_{i}^{*}$. In order to reproduce this from the power spectrum, we need $\mathscr{P}^{*}_{k}$ up to second order in $\varepsilon_{i}^{*}$, which is at the next-to-next-to-leading order (N2LO) for the power spectrum because $\mathscr{P}^{*}_{k}$ at LO is 1. In the late-time limit, the amplitude $\mathscr{P}^{*}_{k}$ up to N2LO with any pivot scale $\tau^{*}$ is given by [17, 18, 19, 20, 4]
$\mathscr{P}^{*}_{k}=\mathcal{A}+\mathcal{B}\ln(-k\tau^{*})+\mathcal{C}\big{(}\ln(-k\tau^{*})\big{)}^{2}+\mathcal{O}\big{(}\varepsilon^{*3},\tau\big{)},$ (4.4)
@@
$\begin{split}&\mathcal{A}=1-2(\lambda+1)\varepsilon_{1}^{*}-\lambda\varepsilon_{2}^{*}+\frac{1}{2}\big{(}\pi^{2}+4\lambda^{2}+4\lambda-6\big{)}\varepsilon_{1}^{*2}\\ &\quad+\frac{1}{12}\big{(}7\pi^{2}+12\lambda^{2}-12\lambda-72\big{)}\varepsilon_{1}^{*}\varepsilon_{2}^{*}+\frac{1}{8}\big{(}\pi^{2}+4\lambda^{2}-8\big{)}\varepsilon_{2}^{*2}+\frac{1}{24}\big{(}\pi^{2}-12\lambda^{2}\big{)}\varepsilon_{2}^{*}\varepsilon_{3}^{*},\end{split}$ (4.5)
$\begin{split}&\mathcal{B}=-2\varepsilon_{1}^{*}-\varepsilon_{2}^{*}+(4\lambda+2)\varepsilon_{1}^{*2}+(2\lambda-1)\varepsilon_{1}^{*}\varepsilon_{2}^{*}+\lambda\big{(}\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\big{)},\end{split}$ (4.6)
$\begin{split}&\mathcal{C}=2\varepsilon_{1}^{*2}+\varepsilon_{1}^{*}\varepsilon_{2}^{*}+\frac{1}{2}\varepsilon_{2}^{*2}-\frac{1}{2}\varepsilon_{2}^{*}\varepsilon_{3}^{*},\end{split}$ (4.7)
@@ 4.2 Consistency relation at NLO
Finally, computing the right hand side of (1) by using (4.3) with (4.4), we obtain
$\begin{split}-\mathcal{P}_{k_{1}}\left(3+k_{2}\frac{d}{dk_{2}}\right)\mathcal{P}_{k_{2}}&=\frac{H^{*4}}{16k_{1}^{3}k_{2}^{3}\varepsilon_{1}^{*2}}\Bigg{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}+2\varepsilon_{1}^{*2}+5\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\\ &\qquad~{}-\big{(}8\varepsilon_{1}^{*2}+6\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\big{)}\big{[}\gamma_{E}-1+\ln(-2k_{2}\tau^{*})\big{]}\\ &\qquad~{}-\big{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}\big{)}^{2}\ln\left(\frac{k_{1}}{k_{2}}\right)\Bigg{)}+\mathcal{O}\big{(}\varepsilon^{*3},\tau\big{)},\end{split}$ (4.8)
which, under $k=k_{2}=k_{3}+\mathcal{O}(k_{1})$, is equal to the bispectrum in the squeezed limit (4.1) with (4.2), thereby verifying the consistency relation at NLO. Note that this has been verified for any choice of the pivot scale $\tau^{*}$.